A 5000 kg elephant is sitting on the jungle floor What is the potential energy of the elephant
1
Expert's answer
2016-01-19T08:28:41-0500
Potential energy value depends on the zero level chosen. Thus, if jungle floor is assumed to be zero level, thus the elephant's PE is zero. However, in case when the center of the Earth is considered to be the reference point, the elephant's PE can be approximated as PE = mgh, where m - is the mass of the elephant; g = 9.8 m/s^2; h - is the distance from the jungle floor to the Earth center (approx. = Earth radius, 6,371 km). Thus,
PE = 5000*9,8*6,371,000 = 3,12*10^11 J

Answer:
elephant's PE related to the jungle floor is zero;
elephant's PE related to the Earth center is 3,12*10^11 J